# Probabilistic Modeling of Antibody Kinetics Post Infection and Vaccination: A Markov Chain Approach

### v1

@@ 8.7 Implications for immunologists
Funding
This research was funded in part by 4-VA, a collaborative partnership for advancing the Commonwealth of Virginia (RAL). P.B. was funded through the NIST PREP, USA grant 70NANB18H162. J.A.W. was funded through NIH grants R21 AI 160334 and R56 AI 178669. J.M.W. was funded through University of Virginia Manning COVID-19 Research Fund. P.S.P. was funded through NIH NIAID R01AI173194. The aforementioned funders had no role in study design, data analysis, decision to publish, or preparation of the manuscript.
Acknowledgements
@@
This work is a contribution of the National Institute of Standards and Technology, USA and is not subject to copyright in the United States. Use of data in this manuscript has been approved by the NIST Research Protections Office under study number ITL-2025-0431. The authors wish to thank Bradley Alpert, Melinda Kleczynski, and Amanda Pertzborn for useful feedback during preparation of this manuscript.
@@ Appendix A Additional time-inhomogeneous two event model details
$\begin{split}&S(4)S(3)S(2)S(2)S(1)S(0)\bm{e}_{1}\\ &=\begin{bmatrix}s_{N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{I}N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{V}N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{I}N}(3)s_{N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{V}N}(3)s_{N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{V}\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{V}\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]+s_{\mathcal{I}\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)\\ s_{\mathcal{V}\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]+s_{\mathcal{V}\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)\\ s_{\mathcal{I}\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]+s_{\mathcal{I}\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)\\ s_{\mathcal{V}\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]+s_{\mathcal{V}\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)\end{bmatrix}\end{split}$ (A1)
We interpret a few entries of (A1) as follows. The first entry, Prob$(X_{4}=N)$, computes the probability that one stays naïve through time step 5; the second entry, Prob$(X_{4}=\mathcal{I})$, gives the probability that one did not become infected on time steps 1, 2, 3, or 4, but did become infected on time step 5. The fourth entry, Prob$(X_{4}=\mathcal{I}^{\prime})$, is the probability that one did not become infected on time steps 1, 2, or 3, but did become infected on time step 4 and is considered previously infected on time step 5 plus the probability that one became infected on a previous time step (computed by adding the individual event probabilities together since they are disjoint), was then considered previously infected, and remained previously infected on subsequent time steps including time step 5. The sixth entry, Prob$(X_{4}=\mathcal{I}\mathcal{I}^{\prime})$, is the probability that one became infected on time steps 1, 2, or 3, then remained previously infected, then becomes newly infected on the fifth time step (indicated by the term $s_{\mathcal{I}\mathcal{I}^{\prime}}(4)$). The quantity in curly braces appears in the fourth entry, too, denoting infection before time step 4. The tenth entry, Prob$(X_{4}=\mathcal{I}^{\prime}\mathcal{I}^{\prime})$, is the probability that one became infected on time step 1 or 2 and becomes newly reinfected on time step 4 and is considered previously reinfected on time step 5, plus the probability that one became infected on time step 1, reinfected at the next possible opportunity (time step 3), and then is considered previously infected thereafter. We confirm that all possible trajectories to $N,\mathcal{I},\mathcal{I}^{\prime},\mathcal{I}\mathcal{I}^{\prime},$ and $\mathcal{I}^{\prime}\mathcal{I}^{\prime}$ on time step 5 are considered. Parallel interpretations hold for quantities involving $\mathcal{V}$,$\mathcal{V}^{\prime}$, and cross events of infection after vaccination or vice versa.
@@
Appendix B Additional data visualizations and models
@@
Figure B1: Log-transformed antibody measurements from the naïve, infected (I), and infected then vaccinated (VI) populations from Dataset 1 with corresponding probability models.
@@
Figure B2: Log-transformed antibody measurements from the vaccinated (V) and boosted (VV) populations from Dataset 1 with corresponding probability models.
@@
Figure B3: Log-transformed antibody measurements from the naïve (synthetic), infected (I), vaccinated (V), and infected then vaccinated (VI) populations from Dataset 2 with corresponding probability models. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level.

### v2

@@ 8.6 Implications for immunologists
9 Relevant mathematical background
The central thesis of our work is that an antibody measurement’s interpretation relies on when the sample was obtained, the prior individual immune events, and the prevalence of events in the population at the time. To that end, this work employs probabilistic modeling to analyze personal antibody kinetics post-immune events such as infection and vaccination. In parallel, we apply time-inhomogeneous Markov chains to study population-level incidence rates and prevalence. Figure 13 shows a schematic that summarizes how the main mathematical ideas of our work fit together. An important accomplishment of this work is the simultaneous tracking of antibody response and immune state across multiple immune events.
Figure 13: Schematic summarizing the main mathematical ideas in this paper.
Personal response models (Figure 13, left).
We first construct probabilistic models of antibody kinetics; such models account for the variation of antibody measurements across individuals. We use the gamma probability density; refer to the naïve model in Figure 3a, for instance. Along the $x$-axis, we have plotted log-scaled antibody levels, while the $y$-axis denotes their relative abundance in the population.
When the disease persists in the population, individuals face multiple immune events, such as reinfection, booster vaccination, or combinations like vaccination followed by infection and vice versa. We model the antibody kinetics of such immune events sequences by considering factors such as the time between immune events. The addition of a new dimension–time since immune event–necessitates a new model depiction. For the models that characterize responses to infection, vaccination, or repeated exposures (e.g., the rest of Figure 3), each vertical slice in time $t$ represents the probability distribution across antibody levels. The $x$-axis denotes the days since the immune event, the $y$-axis denotes the antibody measurement, and the contour lines denote the relative abundance in the population. We strongly encourage readers to watch the videos (described in Appendix E) to visualize of our results. The models are fit to real immune event antibody data via Maximum Likelihood Estimation (MLE), a statistical technique that estimates the parameters associated with a probability distribution that maximizes the likelihood of observing the given data.
Transition probabilities (Figure 13, right).
At the population level, we model how infections and vaccinations spread and accumulate using a time-inhomogeneous Markov chain (see (23), (24), (25)). We specifically study emergent infectious diseases, meaning that immunity emerges de novo, unconfounded by earlier or undocumented immune events. Over time, a fraction of the population moves to the acutely/newly infected ($\mathcal{I}$) state with some time-dependent incidence rate. This fraction then becomes previously infected ($\mathcal{I}^{\prime}$), where an individual possesses time-varying antibody protection that peaks and then wanes. Once a vaccine is developed, analogous states exist for this immune event ($\mathcal{V}$, $\mathcal{V}^{\prime}$). In particular, (24) is a matrix whose entries give the probability of transition from one state to another at a given time $T$. Markov chains are a reasonable approach to track how individuals transition between immune states because the probability of an immune event in the upcoming time step depends on the current immune state. Time-inhomogeneity refers to the fact that the incidence rates and prevalences of these immune events change over time.
Absolute timeline models (Figure 13, bottom).
We combine the personal kinetics and population level changes to model the conditional probability of an antibody measurement. Antibody measurements obtained at a given time from a random sample of the population can be explained as a convolution of the effects of personal antibody kinetics and the population-level prevalences. This defines the probability of obtaining an antibody measurement at a particular time given that the individual faced a certain sequence of immune events so far (see Section 4.2 and Section 5.2).
9.1 Further reading
  • An Introduction to Mathematical Statistics and Its Applications, Larsen and Marx (conditional probability, the gamma distribution, maximum likelihood estimation)
  • Finite Markov Chains and Algorithmic Applications, Häggström (Markov chains, including time-inhomogeneous Markov chains)
  • Plotkin’s Vaccines, Orenstein, Offit, Edwards, & Plotkin (vaccine history, immunology, technology)
  • SARS-CoV-2-infection- and vaccine-induced antibody responses are long lasting with an initial waning phase followed by a stabilization phase, Srivastava et al. Immunity, 57 (2024) pp. 587 - 599 (characteristics of antibody kinetics, hybrid immunity of infection and vaccination)
Declarations
Authorship and contributorship.
All authors have made substantial intellectual contributions to the study conception, execution, and design of the work. All authors have read, reviewed, edited, and approved the final manuscript. In addition, the following contributions occurred: Rayanne A. Luke: Conceptualization, Methodology, Software, Formal analysis, Writing - original draft preparation, Visualization, Funding acquisition. Prajakta Bedekar: Conceptualization, Methodology, Software, Formal analysis, Writing - original draft preparation, Visualization. Lyndsey M. Muehling: Investigation, Data curation. Glenda Canderan: Investigation, Data curation. Yesun Lee: Investigation, Data curation. Wesley Cheng: Investigation, Data curation. Judith A. Woodfolk: Investigation, Resources, Funding acquisition. Jeffrey M. Wilson: Investigation, Resources, Funding acquisition. Pia S. Pannaraj: Conceptualization, Investigation, Resources, Funding acquisition. Anthony J. Kearsley: Conceptualization, Methodology, Supervision, Funding acquisition.
Conflicts of interest.
The authors declare there are no conflicts of interest.
Acknowledgements.
@@
This work is a contribution of the National Institute of Standards and Technology, USA and is not subject to copyright in the United States. Use of data in this manuscript has been approved by the NIST Research Protections Office (study number ITL-2025-0431). The authors thank Bradley Alpert, Melinda Kleczynski, and Amanda Pertzborn for useful feedback during preparation of this manuscript.
Data & code availability.
The authors plan to make code available on GitHub once our paper is accepted for publication. Original data are from Keshavarz et al. [22], Canderan et al. [23], and Congrave-Wilson et al. [24]; see their data availability statements for how to request or access relevant data.
Funding.
R.A.L. was funded by 4-VA, a collaborative partnership for advancing the Commonwealth of Virginia. P.B. was funded through the NIST PREP, USA grant 70NANB18H162. J.A.W. was funded through NIH R21 AI 160334 and R56 AI 178669. J.M.W. was funded through the University of Virginia Manning COVID-19 Research Fund. P.S.P. was funded through NIH NIAID R01AI173194. The aforementioned funders had no role in study design, data analysis, decision to publish, or manuscript preparation.
@@ Appendix A Additional time-inhomogeneous two event model details
$\begin{split}&S(4)S(3)S(2)S(2)S(1)S(0)\bm{e}_{1}\\ &=\begin{bmatrix}s_{N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{I}N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{V}N}(4)s_{N}(3)s_{N}(2)s_{N}(1)s_{N}(0)\\ s_{\mathcal{I}N}(3)s_{N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{V}N}(3)s_{N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{V}\mathcal{I}^{\prime}}(4)\left\{s_{\mathcal{I}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{V}\mathcal{V}^{\prime}}(4)\left\{s_{\mathcal{V}N}(2)s_{N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]\right\}\\ s_{\mathcal{I}\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]+s_{\mathcal{I}\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)\\ s_{\mathcal{V}\mathcal{I}^{\prime}}(3)[s_{\mathcal{I}N}(1)s_{N}(0)+s_{\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)]+s_{\mathcal{V}\mathcal{I}^{\prime}}(2)s_{\mathcal{I}N}(0)\\ s_{\mathcal{I}\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]+s_{\mathcal{I}\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)\\ s_{\mathcal{V}\mathcal{V}^{\prime}}(3)[s_{\mathcal{V}N}(1)s_{N}(0)+s_{\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)]+s_{\mathcal{V}\mathcal{V}^{\prime}}(2)s_{\mathcal{V}N}(0)\end{bmatrix}.\end{split}$ (A1)
[…]
@@
Appendix B Additional population simulation example
$S_{3}={{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}}}{{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}}}{{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}{}{}{}{}{}{}}{{}{}{}{}}}\begin{bmatrix}0.8&0&0&0.01&0.01\\ 0.1&0&0&0.08&0.05\\ 0.1&0&0&0.05&0.06\\ 0&1&0&0.86&0\\ 0&0&1&0&0.88\end{bmatrix}$ (B1)
As expected, the number of vaccinations per individual and stay in previously vaccinated states rises for $S_{3}$ as compared to $S_{2}$; see Figure B1a. Moreover, the time spent in naïve or naïve-like states decreases. We see that individuals frequently move from a previously infected/vaccinated state to another immune event; see Figure B1b.
Figure B1: Simulations for a population with 100010001000 individuals with a time-homogeneous transition matrix $S_{3}$, per 505050 time steps, i.e., approximately two year period. (a) Histograms of counts of total infections, vaccinations, previously infected status, previously vaccinated status, and naïve status. Five representative sequences Light yellow denotes the naïve state (1, $N$), red and pink the newly and previously infected states (2, $\mathcal{I}$ and 4, $\mathcal{I}^{\prime}$), and blue and light blue the newly and previously vaccinated states (3, $\mathcal{V}$ and 5, $\mathcal{V}^{\prime}$).
Appendix C Additional personal trajectories of interest
@@
Figure C1: Log-transformed antibody measurements from the infected then vaccinated (VI) population from Dataset 1 with personal trajectories of interest with corresponding probability models. The relative date of vaccination is marked with a vertical, white dashed line.
In Figure C1 we plot individualized two-event models for vaccination after infection for three subjects of interest. Although the models shown in Section 6 take into account the time between an event and the day of measurement, it would be difficult to visualize this time span in the two-event figures, as there is often no associated measurement value for the day of the event and the extra information may crowd already densely packed data points. When studying personal trajectories, one option is to mark the line connecting available measurements with an additional marker symbolizing the day of the event. Here, while studying personal trajectories one by one, we mark the date of vaccination (after infection or as a booster dose) as a vertical, dashed line. Subject D1-12 has a representative relative time between infection and vaccination (182 days), and exhibits a standard increase in antibody response after vaccination. Both measurement values fall on the higher side of observed values. Subject D1-13 exhibits a strong initial response to infection and a typical decay. Nearly 400 days elapse between their infection and vaccination; likely as a result, they show a strong and sustained response to vaccination. Subject D1-14 is an outlier for their infection response, displaying a very low antibody level almost 200 days after infection, but perhaps in response, they exhibit a very strong increase in antibody level after vaccination.
Appendix D When to measure the event of vaccination?
Because some vaccines are given as a multi-dose sequence, and an individual is not considered fully vaccinated until the sequence is finished, it is interesting to study the effect on our models of setting $t=0$ to represent the day of the first dose of vaccination, rather than the second, as we have done in Section 6. The relative days between the first and second dose of vaccination were available for Dataset 1. We redid our models using dose 1 as $t=0$ (call this Vax Model Dose 1); the results are shown in Table D1 and Figure D1. We note that this allows for more data to be used; in Section 6, only measurements post dose 2 are relevant for the model (call this Vax Model Dose 2; see Figure 3a).
Compared to Vax Model Dose 2, $\theta_{\mathcal{V}_{1}}$ and $\phi_{\mathcal{V}_{1}}$ are an order of magnitude smaller for Vax Model Dose 1; as a result, the peak and decay of Vax Model Dose 1 are smaller and start later, around 60 days as opposed to 10. Ultimately, we are observing the same peak in the two models, which is the response to the second immunization. More of the increase in antibody response is visible when including data after only the first dose in the model; this is paralleled by the infection model (see Figure 3a). Perhaps reassuringly, this modeling choice for the first vaccination event has minimal effect on the model for the booster dose; the parameter values are relatively similar for Vax Model Dose 1 and Vax Model Dose 2.
@@ Appendix D When to measure the event of vaccination?
Figure D1: Log-transformed antibody measurements from the vaccinated (V) and boosted (VV) populations from Dataset 1 with corresponding probability models.
Vaccination 44.1 2.17 2.56
Vaccination,
booster
44.1 2.17 2.56 155 12.6 1.51
Table D1: Optimal parameters from maximum likelihood estimation for Dataset 1 for measurements corresponding to vaccination as the first event. Grayed out numbers indicate variables that are part of the model but are not part of the optimization for that data. Time is measured since first vaccination dose.
Appendix E Supplementary figure captions
Supplementary Figure 1 (GIF).
Log-transformed antibody measurements from the vaccinated then boosted (VV) population from Dataset 1 with personal trajectories of interest with corresponding probability models. The relative date of the booster dose is marked with a vertical, white dashed line. See Figure 11 for accompanying static example images.
Supplementary Figure 2 (GIF).
@@ Appendix E Supplementary figure captions
Log-transformed antibody measurements from the infected, then vaccinated (IV) population from Dataset 2 with personal trajectories of interest with corresponding probability models. The relative date of the vaccination dose is marked with a vertical, white dashed line. Measurements with recorded monotonically decreasing OD values have been AUC-extrapolated using the next titration level. See Figure 12 for accompanying static example images.